The USA won 103 medals. The number of silver medals earned was 4 more than the gold medals. The number of bronze medals was 6 less than the number of gold. How many of each kind of medals did the USA earn?
Answer by haileytucki(390)

x+(x+4)+(x-6)=103
Remove the parentheses that are not needed from the expression.
x+x+4+x-6=103
Since x and x are like terms, add x to x to get 2x.
2x+4+x-6=103
Since 2x and x are like terms, add x to 2x to get 3x.
3x+4-6=103
Subtract 6 from 4 to get -2.
3x-2=103
Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides.
3x=2+103
Add 103 to 2 to get 105.
3x=105
Divide each term in the equation by 3.
(3x)/(3)=(105)/(3)
Simplify the left-hand side of the equation by canceling the common factors.
x=(105)/(3)
Simplify the right-hand side of the equation by simplifying each term.
x=35 gold
35-6= 29 bronze
35+4= 39 silver


Proof:35+29+39
Add 29 to 35 to get 64.
64+39
Add 39 to 64 to get 103.
103